# Fourier Boundary Features Network with Wider Catchers for Glass Segmentation

### v1

@@ V Conclusion and Future Work
At the same time, we also objectively noted that the proposed method still has some limitations. When segmenting specific objects, as shown in Figure 11, for buildings covered by large glass groups, the network cannot effectively explore such boundaries to achieve fine and complete segmentation. Similarly, when the stained glass with decorative function is segmented, the method is susceptible to interference from the glass itself, so that the area where the glass is located cannot be accurately identified. In the future, our work will explore the application of shallow network architecture in more scenarios like buildings with large areas of glass or special glass.
@@ Acknowledgment
This work was supported in part by the Natural Science Foundation of Sichuan (24NSFJQ0069), Science and Technology Service Network Initiative (KFJ-STS-QYZD-2021-21-001), the Talents by Sichuan provincial Party Committee Organization Department, and Chengdu - Chinese Academy of Sciences Science and Technology Cooperation Fund Project (Major Scientific and Technological Innovation Projects).

### v2

@@ V Conclusion and Future Work
At the same time, we also objectively noted that the proposed method still has some limitations. When segmenting specific objects, as shown in Figure 14, for buildings covered by large glass groups, the network cannot effectively explore such boundaries to achieve fine and complete segmentation. When the stained glass with decorative function is segmented, the results of the method are not ideal.
After analysis, we believe that the possible reasons are as follows: Different from conventional glass, colorful decorative glass to improve the decoration, the color fullness of the glass itself and the complexity of the pattern will be greatly increased, which leads to an increase in the difficulty of retaining the general characteristics of glass, such as transparency, reflective performance, and so on. From the perspective of semantic information, the semantic features of such glass are more difficult to distinguish from context. Furthermore, decorative colored glass with high texture and color richness poses challenges for extracting boundary information. And the proposed boundary constraint method requires the support of real data, thus placing higher demands on the quality of data annotation, and is affected by the subjectivity and accuracy of manual annotation.
Therefore, in future work, it is necessary to introduce more image data containing complex boundaries or colored glass, so that the model can better complete diverse glass image segmentation tasks. Meanwhile, we will explore semi-supervised or unsupervised learning methods to reduce reliance on high-quality annotated data.
@@ Acknowledgment
This research was partly supported by the Sichuan Science and Technology Program (2024NSFJQ0035), the Talents by Sichuan provincial Party Committee Organization Department, and Chengdu - Chinese Academy of Sciences Science and Technology Cooperation Fund Project (Major Scientific and Technological Innovation Projects).